OpenStudy (firejay5):

For each problem, you must: 1. Define the variables 2. Write the function to be maximized or minimized. 3. Write a system of inequalities. 4. Graph and write the vertices of the feasible region. 5. Solve the problem and answer the question. The Conway High Band is selling caramel and crunch bars to raise money for their trip to Florida in 2012. The caramel bars are packaged 20 bars per case. They pay $30 per case of caramel and make a profit of $17 per case. The crunch bars are packaged 24 bars per case and CHS pay $26 per case and makes a profit of $15 per case. They order no more than 3,000 caramel and crunch bars together and spend nor more than $4,000. How many cases of caramel and crunch bars would Mr. Cunningham need to buy to maximize their profits?

OpenStudy (goformit100):

i am weak at it

hero (hero):

No, never do that. Only post one question at a time. You can close this question now and post a new open question that I'm already here if you want.

OpenStudy (firejay5):

can you view 2 questions at a time

hero (hero):

You can close this question and open another question. I will work on this one. Someone else can at least look at the other one. If no one else helps with the other question, I can help with that too.

OpenStudy (firejay5):

I need to have this done by today at least

hero (hero):

1. Define Variables: x = no of cases of caramel bars y = no of cases of crunch bars 2. Write Associated Equations and Inequalities: (Amount) 20x + 24y <= 3000 (Cost) 30x + 26y <= 4000 Profit) 17x + 15y = P(x,y) 3. Simplify each inequality and solve the system: 20x + 24y <= 3000 4(5x + 6y) <= 4(750) 5x + 6y <= 750 30x + 26y <= 4000 2(15x + 13y) <= 2(2000) 15x + 13y <= 2000 Write inequalities as a system: 5x + 6y <= 750 15x + 13y <= 2000 Multiply the first inequality by 3: 15x + 18y <= 2250 15x + 13y <=2000 Isolate 15x in each inequality: 15x <= 2250 - 18y 15x <= 2000 - 13y Subtract the inequalities: 0 <= 250 - 5y 5y <= 250 y <= 250/5 y < = 50 Find x by substituting y back in: 15x + 13(50) <=2000 15x + 650 <= 2000 15 x <= 2000 - 650 15x <= 1350 x <= 90 Therefore the max values are x = 90, y = 50, meaning 90 caramel bar cases were bought and sold and 50 crunch bar cases were bought and sold 4: Use the Profit Equation to find the Total Profit: P(x,y) = 17x + 15y P(x,y) = 17(90) + 15(50) P(x,y) = 2280 Therefore the total profit is $2280 5. Find the Total Cost using the Cost Function: C(x,y) = 30x + 26y C(x,y) = 30(90) + 26(50) C(x,y) = 4000 Therefore the total cost was $4000, the maximum that could be spent.

OpenStudy (firejay5):

@Hero What's #3

hero (hero):

I wrote the solution to problem. It's up to you to decipher it.

OpenStudy (firejay5):

so are there 3 functions to be maximized @Hero

hero (hero):

Not exactly. System of inequalities was used to find the maximum values of the variables.

OpenStudy (firejay5):

what's the function to be maximized

hero (hero):

You want to maximize profit of course.

hero (hero):

I cannot help you with every single category bro. Some things you have to be able to figure out yourself.

OpenStudy (firejay5):

aww really

hero (hero):

Sadly, yes. You need to be able to use your own thinking ability to figure out how to solve problems. Its unfortunate, but it is a reality of life. If you don't acquire the ability to do this, you may have some tough challenges facing you in life.

OpenStudy (firejay5):

poopy

OpenStudy (firejay5):

@Hero Where did you get 90 and 50